The following companies are trying to decide whether to change their accounts receivables collection ?? ach company Three Rivers Industries Twin Peaks Ltd. Single Signals Inc. Current receivables policy Unit sale price Contribution margin per unit Units sold per year on credit Average collection period in days Bad debts per year nterest rate on line of credit ts cash flow e controller $125 $ 50 120,000 72 $175,000 11% $250 75,000 $149,000 $550 $140 25,000 100 $160,000 9% 7% Proposed new receivables policy Revised average collection period in days Cost % reduction in unit sales 70 30 $45,000 7% $60,000 $89,000 10% Required: Compute the cost of the current receivables policy for each company 8.8 Required: lie the data in Problem and Case 8.7 to compute the cost of switching to a new accounts receivables policy for ach company
Answer #1

1. At current policy, the total cost will be as follows :

Three Rivers Industries Twin Peaks Ltd. Single Signal Inc. Total A. Unit sale price B. Contribution margin per case C. 125 50

Note 1 - average collection period is 72 days in first column and there are 360 days in a year which means that the cash was collected five times in a year (360/72 = 5 times). - this is how the other 2 calculated.

Note 2 - G (Collection in single time) - We have collected 5 times and total credit sale is 15,000,000 i.e. in single time we hve collected 15,000,000/5 = 3,000,000. - this is how the other 2 calculated.

Note 3 - As 3,000,000 was collected 5 times in a year the interest cost will be calculated 5 times for 72 days in one portion. (3,000,000x11%x72/360)*5 = 330,000 - this is how the other 2 calculated.

Note 4 - we have incurred interest cost and bad debts cost - the total of these will be the total cost of receivables.

2. Cost of the new policy-

Three Rivers Industries Twin Peaks Ltd. Single Signal Inc. Total A. Unit sale price B. Contribution margin per case C. Units

The total cost has come down to 618,653 because of the lower interst rates and low bad debts costs.

thus the new policy is beneficial if we can work over these number practically.

THanks